# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

### v1

@@ B.3. Proof of Proposition 7
To derive the result of , we introduce two lemmas. First, we establish that the difference in the expectation of certain functions can be upper bounded by moments of their inputs.
@@
Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Let $\mathbf{X}$ and $\mathbf{W}$ denote matrices $(X^{\top}_{1},\cdots,X^{\top}_{n})^{\top}$ and $(W^{\top}_{1},\cdots,W^{\top}_{n})^{\top}$, respectively. Then, there exists some $C_{0}=C_{0}(\tau,q)>0$ such that for any finite set $\mathcal{W}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, any $\rho\in(0,1)$, and any $g\in C^{3}(\mathbb{R}^{n\times d})$ satisfying $\max_{i\in[n],j\in[p]}|E[\partial_{i,j}g(\mathbf{Z}^{0}_{i})]|<\infty$ and $\max_{i\in[n],j,k\in[p]}|E[\partial_{i,j}\partial_{i,k}g(\mathbf{Z}^{0}_{i})]|<\infty$, we have
@@
×$\times\left(\frac{1}{n}\sum_{i=1}^{n}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\right.$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$\left.+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right).$
where we set $\tau\leq\frac{n}{d}\leq\frac{1}{\tau}$ and denote $C^{j}_{i\ell}$ as the $(i,\ell)$ element of $\mathbf{Z}_{i,j}^{0}$. The summation of the last term expresses the sum of all the combinations of $\mathbf{{\bar{q}}}_{0}$ random variables exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$.
@@
$\lesssim_{\tau,q}\frac{d^{(\bar{\mathbf{q}}_{0}/2)+1}}{n}\max\{\max_{1\leq k\leq p}(\|X_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})},\max_{1\leq k\leq p}(\|W_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})}\}+(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|}))^{(\bar{\mathbf{q}}_{0}/2)},$
@@
From Lemma 13 and 14, we have
@@
×$\times\left\{\frac{1}{n}\sum_{i=1}^{n}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\right.$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$+\left.\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C_{ik_{1}}\cdots C_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right\}.$
@@
$E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]=$ $E\left[\sum_{\ell=1}^{|B_{j}|}\sum_{\ell^{\prime}=1}^{|B_{j}|}\sum_{\ell^{\prime\prime}=1}^{|B_{j}|}|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}|\right]$
$\leq$ $|B_{j}|^{3}\max_{\ell,\ell^{\prime},\ell^{\prime\prime}\in[|B_{j}|]}E\left[|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}|\right]$
By the Cauchy-Schwartz inequality, for any ,,′′$\ell,\ell^{\prime},\ell^{\prime\prime}\in[|B_{j}|]$, we have
$E\left[|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}|\right]$ $\leq\|(X^{(B_{j})}_{i})_{\ell}(X^{(B_{j})}_{i})_{\ell^{\prime}}\|_{2}\|(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}\|_{2}$
$\leq\|(X^{(B_{j})}_{i})_{\ell}\|_{4}\|(X^{(B_{j})}_{i})_{\ell^{\prime}}\|_{4}\|(X^{(B_{j})}_{i})_{\ell^{\prime\prime}}\|_{4}$
$\leq\max_{\ell\in[|B_{j}|]}\|(X^{(B_{j})}_{i})_{\ell}\|^{3}_{4}$
@@
$E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(X^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]\leq|B_{j}|^{\bar{\mathbf{q}}_{0}+3}\max_{\ell\in[|B_{j}|]}\|(X^{(B_{j})}_{i})_{\ell}\|^{\bar{\mathbf{q}}_{0}+3}_{2^{(\bar{\mathbf{q}}_{0}+4)/2}}.$
Then, from Lemma 6 and 14, we obtain
$\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]$ $\lesssim\frac{pd^{2}}{n^{3/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\},$
@@
$\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
@@
$\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}\sum_{j=1}^{k}E\left[\left(\sum_{\ell=1}^{|B_{j}|}\left|(Z^{(B_{j})}_{i})_{\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C_{ik_{1}}\cdots C_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
$\lesssim_{\tau,q}$ $\frac{pd^{3+(\bar{\mathbf{q}}_{0}/2)}}{n^{5/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\}\left(\max\{\max_{1\leq k\leq p}\|X_{i,k}\|^{(\bar{\mathbf{q}}_{0})}_{2^{(\bar{\mathbf{q}}_{0}/2)}},\max_{1\leq k\leq p}\|W_{i,k}\|^{(\bar{\mathbf{q}}_{0})}_{2^{(\bar{\mathbf{q}}_{0}/2)}}\}\right)$
$+$ $\frac{pd^{2}}{n^{3/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\}\left(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|})\right)^{\bar{\mathbf{q}}_{0}/2}.$
@@ B.4. Proof of Lemma 8 and Lemma 9
$\leq 2\mathcal{D}_{\ell}(\rho)\delta L^{q_{0}q_{1}}_{n}Av(\{L^{q_{1}}_{\ell_{i}}\})\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)$

### v2

@@ B.3. Proof of Proposition 7
To derive the result , we introduce two lemmas. First, we establish that the difference in the expectation of certain functions can be upper bounded by moments of their inputs.
@@
Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Let $\mathbf{X}$ and $\mathbf{W}$ denote matrices $(X^{\top}_{1},\cdots,X^{\top}_{n})^{\top}$ and $(W^{\top}_{1},\cdots,W^{\top}_{n})^{\top}$, respectively. Then, there exists some $C_{0}=C_{0}(\tau,q)>0$ such that for any finite set $\mathcal{W}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, any $\rho\in(0,1)$, and any $g\in C^{3}(\mathbb{R}^{n\times d})$ satisfying $\max_{i\in[n],m\in[p]}|E[\partial_{i,m}g(\mathbf{Z}^{0}_{i,j})]|<\infty$ and $\max_{i\in[n],m,m^{\prime}\in[p]}|E[\partial_{i,m}\partial_{i,m^{\prime}}g(\mathbf{Z}^{0}_{i,j})]|<\infty$, we have
@@
$\times\left(\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]\right$.
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$\left.+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right).$
where we set $\tau\leq\frac{n}{p}\leq\frac{1}{\tau}$ and denote $C^{j}_{i\ell}$ as the $(i,\ell)$ element of $\mathbf{Z}_{i,j}^{0}$. The summation of the last term expresses the sum of all the combinations of $\mathbf{{\bar{q}}}_{0}$ random variables exerted from the $i$-th row of $\mathbf{Z}_{i,j}^{0}$ except those in $B_{j}$.
@@
$\lesssim_{\tau,q}\frac{d^{(\bar{\mathbf{q}}_{0}/2)+1}}{n}\max\left\{\max_{1\leq k\leq p}(\|X_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})},\max_{1\leq k\leq p}(\|W_{i,k}\|_{2^{(\bar{\mathbf{q}}_{0}/2)}})^{(\bar{\mathbf{q}}_{0})}\right\}$
$\quad+\left(\max_{1\leq i\leq k}\lambda_{\max}(\Sigma_{|B_{i}|})\right)^{(\bar{\mathbf{q}}_{0}/2)},$
@@
From Lemma 15, we have
@@
$\times\left\{\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]\right$.
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]$
$+\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
$+\left.\frac{1}{n}\sum_{i=1}^{n}L^{\mathbf{q}}_{\ell_{i}}\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C^{j}_{ik_{1}}\cdots C^{j}_{ik_{\bar{\mathbf{q}}_{0}}}\right]\right\}.$
@@
$E\left[\left(\sum_{\ell\in B_{j}}\left|X_{i,\ell}\right|\right)^{3}\right]=$ $E\left[\sum_{\ell\in B_{j}}\sum_{\ell\in B_{j}}\sum_{\ell^{\prime\prime}\in B_{j}}|X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}|\right]$
$\leq$ $|B_{j}|^{3}\max_{\ell,\ell^{\prime},\ell^{\prime\prime}\in B_{j}}E\left[|X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}|\right].$
By the Cauchy-Schwartz inequality, for any $\ell,\ell^{\prime},\ell^{\prime\prime}\in B_{j}$, we have
$E\left[|X_{i,\ell}X_{i,\ell^{\prime}}X_{i,\ell^{\prime\prime}}|\right]$ $\leq\|X_{i,\ell}X_{i,\ell^{\prime}}\|_{2}\|X_{i,\ell^{\prime\prime}}\|_{2}$
$\leq\|X_{i,\ell}\|_{4}\|X_{i,\ell^{\prime}}\|_{4}\|X_{i,\ell^{\prime\prime}}\|_{4}$
$\leq\max_{\ell\in B_{j}}\|X_{i,\ell}\|^{3}_{4}.$
@@
$E\left[\left(\sum_{\ell\in B_{j}}\left|X_{i,\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]\leq|B_{j}|^{\bar{\mathbf{q}}_{0}+3}\max_{\ell\in B_{j}}\|X_{i,\ell}\|^{\bar{\mathbf{q}}_{0}+3}_{2^{(\bar{\mathbf{q}}_{0}+4)/2}}.$
Then, from Lemma 6 and 16, we obtain
$\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]$ $\lesssim\frac{pd^{2}}{n^{3/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\},$
@@
$\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{\bar{\mathbf{q}}_{0}+3}\right]$
@@
$\sum_{j=1}^{k}\max_{Z_{i}^{(B_{j})}\in(X_{i}^{(B_{j})},W_{i}^{(B_{j})})}E\left[\left(\sum_{\ell\in B_{j}}\left|Z_{i,\ell}\right|\right)^{3}\right]\sum_{k_{1},\cdots,k_{\bar{\mathbf{q}}_{0}}\notin B_{j}}E\left[C_{ik_{1}}\cdots C_{ik_{\bar{\mathbf{q}}_{0}}}\right]$
$\lesssim_{\tau,q}\frac{pd^{3+(\bar{\mathbf{q}}_{0}/2)}}{n^{5/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\}$
$\qquad\times\left(\max\{\max_{1\leq k\leq p}\|X_{i,k}\|^{(\bar{\mathbf{q}}_{0})}_{2^{(\bar{\mathbf{q}}_{0}/2)}},\max_{1\leq k\leq p}\|W_{i,k}\|^{(\bar{\mathbf{q}}_{0})}_{2^{(\bar{\mathbf{q}}_{0}/2)}}\}\right)$
$+\frac{pd^{2}}{n^{3/2}}\max\left\{\max_{1\leq k\leq p}\|X_{i,k}\|^{3}_{4},\max_{1\leq k\leq p}\|W_{i,k}\|^{3}_{4}\right\}\left(\lambda_{\max}(\Sigma)\right)^{\bar{\mathbf{q}}_{0}/2}.$
@@ B.4. Proof of Lemma 8 and Lemma 9
$\leq 2\mathcal{D}_{\ell}(\rho)\delta L^{q_{0}q_{1}}_{n}Av(\{L^{q_{1}}_{\ell_{i}}\})\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right),$